OpenStudy (anonymous):

your friend deposits $6,000in an investment account that earns 7.1% annual interest. Find the balance after 12 years when the interest is compounded quarterly. I know the formula is a= p (1+ r/n)^nt

OpenStudy (anonymous):

the answers are a. $13,665.48 B. $14,030.72 C.$14,065.98 D.$13,961.26

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@jdoe0001

OpenStudy (qchronicles):

What are you having trouble with?

OpenStudy (qchronicles):

Hello are you there?

OpenStudy (anonymous):

i need to solve the problem

OpenStudy (anonymous):

idk where to plug the numbers tho

OpenStudy (qchronicles):

No I mean you have the formula so what part of it are you having problems with

OpenStudy (qchronicles):

Ok hang on a sec

OpenStudy (jdoe0001):

\(\qquad \textit{Compound Interest Earned Amount}\\ A=P\left(1+\frac{r}{n}\right)^{nt} \qquad \begin{cases} A=\textit{current amount}\\ P=\textit{original amount deposited}\to &\$6,000\\ r=rate\to 7.1\%\to \frac{7.1}{100}\to &0.071\\ n=\textit{times it compounds per year}\to &4\\ t=years\to &12 \end{cases}\)

OpenStudy (qchronicles):

Thx jdoe

OpenStudy (qchronicles):

I think that sums it up

OpenStudy (jdoe0001):

quarterly, means every quarter, there are 4 quarters for a whole, so the cycle or period is 4

OpenStudy (anonymous):

but i did that and i ended up with like 5,000 which was not any of the answer choices

OpenStudy (qchronicles):

Can you show me what you did that way I can see were u went wrong?

OpenStudy (anonymous):

one sec I'm trying it

OpenStudy (qchronicles):

k

OpenStudy (anonymous):

okay i got it right this time thanks :)

OpenStudy (qchronicles):

Ok medal?

OpenStudy (anonymous):

can you help with others

OpenStudy (qchronicles):

how many?

OpenStudy (anonymous):

llike 5 but they are different types

OpenStudy (anonymous):

do you have a snap chat so i can show you the problem?

OpenStudy (qchronicles):

No I dont sorry

OpenStudy (qchronicles):

If u have time to type them Il answer them tho.

OpenStudy (anonymous):

okay i do :)

OpenStudy (qchronicles):

k

OpenStudy (anonymous):

the next one is simplify the expression

OpenStudy (qchronicles):

k

OpenStudy (anonymous):